OpenStudy (anonymous):

For any time t≥0, if the position of a particle in the xy-plane is given by x=t^2+1 and y=ln(2t+3), then the acceleration vector is

OpenStudy (amistre64):

isnt it the derivative of the unit tangent vector?

OpenStudy (amistre64):

r(t) = (x(t), y(t)) r'(t) = (x'(t), y'(t)) , get the unit form of that and derive again ....

OpenStudy (anonymous):

I thought it was a parametrization problem

OpenStudy (amistre64):

parametric equations are just a vector from the origin pointing to the x,y point at its current location

OpenStudy (amistre64):

if your simply trying to get dy/dx, then get (dy/dt)/(dx/dt) and see if its simpler to play with

OpenStudy (anonymous):

so i get x and y and derive it

OpenStudy (amistre64):

y=ln(2t+3) x=t^2+1 y'=2/(2t+3) x'=2t dy/dx = (2t^2+3t)^(-1)

OpenStudy (amistre64):

the key here is this; what we have here is not dy/dx perse so just deriving again is not going to work out so well. what we actually want for acceleration is:\[\frac{\frac{d}{dt}\frac{dy}{dx}}{dx/dt}\] if memory serves

OpenStudy (anonymous):

okay then so, to find the acceleration vector, V(x), one needs to first derive x and y right

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

so it would be in this form, (x', y')

OpenStudy (amistre64):

yep, thats the tangent vector ... we want to turn that into a unit vector and derive again

OpenStudy (anonymous):

thanks i think i can take it from here

OpenStudy (amistre64):

good :)